B spends 3/8 of his pocket money.If he has $5.20 left, how much did he have to begin with?
Answer by rwm(914)

he spent 3/8 so he has 5/8
5/8*x=5.20
x=5.2*8/5
x=8.32
5/8*8.32=5.20